# Hadamard Representations: Augmenting Hyperbolic Tangents in RL

### v3

@@ 3 Augmenting Hyperbolic Tangents
Hyperbolic Tangent:
In the case of the hyperbolic tangent, product saturation only occurs if strictly both neurons are saturated. This results in a probability of $\mathbf{p}\cdot\mathbf{p}=\mathbf{p}^{2}$. Taking a product of hyperbolic tangent activated neurons thus reduces the probability of neuron saturation from $\mathbf{p}$ to $\mathbf{p}^{2}$.
Sigmoid:
Rectified Linear-Unit:
In the case of a ReLU activation, we also assume the probability of a single neuron dying to be $\mathbf{p}$. As we look at the product of two neurons, the probability that one of the two neurons does not saturate is therefore $1-\mathbf{p}$, and the probability that both neurons do not saturate is $(1-\mathbf{p})^{2}$. The probability that at least one neuron saturates is thus equal to $1-(1-\mathbf{p})^{2}=2\mathbf{p}-\mathbf{p}^{2}$. As the ReLU saturation results in strict zeroes, this results in the product also being zero. Taking a product of ReLU activated neurons therefore increases the final neuron saturation probability from $\mathbf{p}$ to $2\mathbf{p}-\mathbf{p}^{2}$. For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.3.
Activation Function Dying Neuron Probability Probability with HR Difference
Hyperbolic Tangent $\mathbf{p}$ $\mathbf{p}^{2}$ $-(\mathbf{p}-\mathbf{p}^{2})$
Sigmoid $\mathbf{p}$ $\mathbf{p}$ 0
ReLU $\mathbf{p}$ $2\mathbf{p}-\mathbf{p}^{2}$ $+(\mathbf{p}-\mathbf{p}^{2})$
4 Experiments
The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017) and the parallelized Q-network (PQN) (Gallici et al., 2024). For more details on the PPO, DQN and PQN implementation, we refer the reader to Appendix A.1. In line with Gulcehre et al. (2022), for our qualitative experiments, a Hadamard Representation is only applied to the last hidden layer $z_{t}\in\mathbb{R}^{512}$. We perform qualitative experiments on 8 common, non-exploration driven Atari games. We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance. For more information on the dying neuron and effective rank calculations, we refer the reader to Appendix. B.1 and Appendix. B.2, respectively. The performance is aggregated over 8 environments and normalized using the ReLU baseline scores. Finally, the PQN experiments are conducted with an HR on all hidden layers, showing the Median Human-Normalized scores on 51 Atari games.
(a) Dead Neurons
(b) Effective Rank
Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames). Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior during training. A Hadamard product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
@@ 4 Experiments
As mentioned in Section 3 and defined in Eq. 1, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$. The amount of dead neurons over time when training on Atari can be seen in Fig.  6(a). For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ die during training, while for the sigmoid and hyperbolic tangent activation this number is around 40%. When using a hyperbolic tangent Hadamard representation, a reduction in dead neurons as compared to using a single hyperbolic tangent can be observed. We credit this to the inherent ability of a Hadamard product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 3. Quantitative results of dead neurons can be found in Table 2, which is correlated with earlier predictions from Table 1.
Table 2: Dying neuron fractions with and without a Hadamard representation (HR).
Activation Function Dead Neuron Fraction Dead Neuron Fraction (HR) Difference
Hyperbolic Tangent 0.39 0.30 -23%
@@
We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 6(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, similar to the results in our supervised learning experiments (see Fig. 2), using a Hadamard representation with hyperbolic tangents significantly improves the effective rank, which is strongly correlated to a network’s ‘expressivity’. Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected.
Analyzing Performance in Atari
@@
The influence of a Hadamard representation on downstream performance can be seen in Fig. 7. Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline. Furthermore, a comparison is made with the novel Rational (Delfosse et al., 2024) activation function as the activation in the final hidden layer. Fig. 7(a) shows that, although the Rational activation seems to be a stable learnable activation, it remains comparable to the ReLU.
(a) Normalized Score
(b) Ablations
@@
Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), the Tanh (HR) significantly outperforms the baseline Tanh. The recently proposed Rational Activation (Delfosse et al., 2024) seems to have comparable performance to the ReLU in the 40M frame setting. In (b), several ablations including layer normalization (Ba et al., 2016) and an increased dimension of $z_{t}\in\mathbb{R}^{1024}$ with a reduced $\alpha=5e^{-5}$ on hyperbolic tangent activations are shown. Note that no hyperparameter changes are done for the Hadamard Representation.
Figure 8: Additional ablations run on DQN in the Atari domain. Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product.
@@
Examining further ablations in Fig. 8 shows that using a sophisticated piece-wise linear function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product. We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$. Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
@@
To evaluate results on a policy-based algorithm, additional experiments are run using the PPO algorithm (Schulman et al., 2017). For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2). The results can be found in Fig. 9. Similarly to the results on DQN, the Tanh (HR) exhibits the highest effective rank and the best performance. The Tanh (HR) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations can be relatively unreliable across algorithms.
(a) Normalized Score
(b) Effective Rank
Figure 9: Normalized performance (a) and effective rank (b) in the Atari domain for 10M iterations (40M Frames) when training the PPO algorithm (Schulman et al., 2017).
@@
To more reliably test the performance gains one can achieve using Hadamard Representations, we evaluate their effects on the Parallelized Q-Network (PQN) in the full Atari suite without hard-exploration environments, for a total of 51 games with 5 seeds per game.
@@
Figure 10: Median Human-Normalized scores on PQN in 51 Atari games for 5 seeds over 40M frames. Labels represent MLP / CNN activations. Employing a Hadamard Representation in the Convolutional layers as well as the MLP layer provides a significant performance improvement over the recent PQN baseline, as well as a CReLU (Abbas et al., 2023) implementation of PQN.
@@
PQN has recently emerged as a vectorized-friendly version of DQN, reporting both a speed up from DQN as well as better convergence (Gallici et al., 2024). We run tests implementing the Hadamard Representation in all hidden layers. The median human-normalized scores after training for 40M frames are shown in Fig. 10. We can clearly see a similar trend for the 51-game Atari suite, where using a Hadamard representation for all layers significantly increases the median performance as compared to the baseline PQN and CReLU (Abbas et al., 2023)
Hyperbolic Tangents turn Weights into Biases
Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations. As also indicated by Teney et al. (2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood. However, in the case of dying neurons, we aim to shed some light on this phenomenon.
Theorem 4.1.
When any set of neurons $\alpha^{j}$ in a hidden layer $z^{j}$ collapses into nonzero values, the output to the next layer effectively changes from ($A^{j}z^{j}+B^{j}$) to ($A_{-*}^{j}z_{i-*}^{j}+B^{j+1}$ + $A_{*}^{j}z_{i*}^{j}$), where $A_{-*}^{j}z_{-*}^{j}$ represent the active neurons multiplied by their corresponding forward-connected weights and $A_{*}^{j}z_{i*}^{j}~{}=~{}B_{*}^{j+1}$ represent the dead neurons, inducing a hidden bias.
Proof.
Take a set of neurons $\alpha^{j}_{i}$ and forward connected weights $\bm{w}^{j}_{\alpha_{i}}$ in layer $z^{j}$. The influence of these neurons on the next layer $z^{j+1}$ is calculated as:
$z^{j+1}=\sum_{i}\alpha^{j}_{i}\bm{w}^{j}_{\alpha_{i}}+B^{j+1}.$ (2)
If the set of neurons dies and collapses into 0 ($\alpha^{j}_{i}=0~{}\forall~{}s\in\mathcal{S}$), which occurs when using ReLU activations, the influence on the next layer becomes 0, representing basic pruning:
$\sum_{i}0\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=0\quad\forall s\in\mathcal{S}.$ (3)
Where $\bm{\bar{w}}^{j}_{\alpha_{i}}$ is the set of weights connected to the dying neurons. However, for a hyperbolic tangent activation, if the set of neurons saturate into either -1 or 1 ($\alpha^{j}_{i}=\{1,-1\}\forall s\in\mathcal{S}$), the output is
$\sum_{i}\{1,-1\}\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=B_{*}^{j+1}\quad\forall s\in\mathcal{S}.$ (4)
As a result, the weight vector $\bm{\bar{w}}^{j}_{\alpha_{i}}$ corresponding to the dying activations only influences a bias $B^{j+1}_{*}$ on the resulting hidden layer. Note that the bias $B^{j+1}_{*}$ is constant for any input observation. ∎
This emergent bias can hinder the optimization process by introducing unintended fixed contributions to a networks’ next hidden layer or Q-values, reducing the flexibility of the network’s representations and potentially reducing its performance in RL or supervised learning.
5 Related Work
Network Capacity in RL.
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues. Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022). Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and Sokar et al. (2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training and Delfosse et al. (2024) applied rational activations (Molina et al., 2019) in RL to increase plasticity. Concurrent work by Dohare et al. (2024) used continual back-propagation to further alleviate plasticity loss. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024). This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
Network Architecture in RL
The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject. Work by Srivastava et al. (2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997). Using these ideas, the Resnet was invented (He et al., 2016) and also showed strong performance in combination with RL (Espeholt et al., 2018). Further work by Henderson et al. (2018) showed differences in RL performances over different network architectures and nonlinear activations. Work by Abbas et al. (2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch. Finally, recent work by Grooten et al. (2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
@@ 6 Conclusions and Discussion
This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem. Based on this analysis, we propose a novel representation architecture, the Hadamard Representation (HR), which enhances an encoder’s final hidden layer by taking the Hadamard product with a parallel, independently parameterized activation layer. We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness. In DQN, PQN and PPO, this approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension. Finally, we have given insights into the ability of the ReLU activation to be less affected the symptoms of dying neurons, as opposed to non-zero saturating activations such as the hyperbolic tangent.
Future work could focus on further identifying the intricacies of the effects that different activation functions have on the resulting representation, in an attempt to push the potential of non piece-wise linear activations in reinforcement learning or even in a supervised learning setting (see Fig. 2). Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al. (2023) or in Delfosse et al. (2024) could be promising.
7 Limitations
Using a Hadamard representation will double the incoming weights connected to a hidden layer. However, in Fig. 10, a comparison against the CReLU is shown, which equally doubles the network’s parameters. Furthermore, recent work shows that a scaling baselines in Atari often leads to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024). Second, due to the vast amount of activation functions and the theoretical sound candidacy for the hyperbolic tangent, we have mostly focused our research on the combination of hyperbolic tangents. Further research into multiplicative representations could find even better activation combinations. A hyperparameter search could also give more insights into the strengths of the Hadamard Representation, since the baselines are specifically tuned for the ReLU activation. Finally, as this research performs experiments on DQN, PPO and PQN, integration of a Hadamard representation into more complex algorithms and architectures such as Rainbow (Hessel et al., 2018), Impala (Espeholt et al., 2018) or SPR Schwarzer et al. (2021) would be interesting.
Reproducibility Statement
Constructing a Hadamard representation is a relatively straightforward process, of which the Pytorch pseudocode can be found in Appendix A.1. The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022). The PQN algorithm uses the implementation in the Github repository of Gallici et al. (2024). Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.

### v4

@@ 5 Experiments
We analyze the effect of a Hadamard representation on a hidden layer’s fraction of dying neurons, its effective rank and the downstream performance. In the qualitative analysis, we evaluate DQN (Mnih et al., 2015) and PPO (Schulman et al., 2017) on 8 common, non-exploration driven Atari environments for 40M frames. Finally, a larger performance analysis is done using the recent Parallelized Q-Network (PQN) (Gallici et al., 2024) on 51 Atari environments for 40M frames.
(a) Dead Neurons (DQN)
(b) Effective Rank (DQN)
(c) Normalized Score (DQN)
Figure 7: (a) The average fraction of dead neurons, (b) the average effective rank (Kumar et al., 2021), and (c) the baseline-normalized score when training DQN in the Atari domain for 10M iterations (40M frames). Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior. A Hadamard product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
@@
As defined in Eq. 1, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$. Over 8 Atari environments, the average amount of dead neurons during training can be seen in Fig.7(a). For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ die during training, while for the sigmoid and hyperbolic tangent activation this number is around 40%. When using a hyperbolic tangent Hadamard representation, a reduction in dead neurons as compared to using a single hyperbolic tangent can be observed. We credit this to the inherent ability of a Hadamard product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 4. Quantitative results of dead neurons can be found in Table 2, which tends to confirm the predictions from Table 1. As compared to the Hadamard representation, layer normalization notably reduces dead dead hyperbolic tangents to an even lower value. Interestingly, the effects of layer normalization show that even if dead neurons are reduced more strongly, it does not always coincide with a higher effective rank (see Fig. 7(b). On the other hand, the Hadamard representation does seem to have a positive effect on both metrics.
Table 2: Average dying neuron fractions in 8 Atari Games with and without a Hadamard representation (HR).
Activation Without HR With HR $\Delta$
Tanh 0.39 0.30 -23%
@@
We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 7(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, similar to the results in our supervised learning experiments (see Fig. 2), using a Hadamard representation with hyperbolic tangents significantly improves said effective rank, which is strongly correlated to a network’s ‘expressivity’. Alternatively, we found that employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected from earlier predictions in Table 1. More details on the ReLU Hadamard representation, dying neuron calculations and effective rank calculations can be found in Appendices C.4, B.1 and B.2, respectively.
Performance in Atari
@@
The influence of a Hadamard representation on downstream performance is visualized in Fig. 7(c). Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline. Furthermore, a comparison is made with the novel Rational (Delfosse et al., 2024) activation function as the activation in the final hidden layer. Fig. 7(c) shows that, although the Rational activation seems to be a stable learnable activation, it remains comparable to the ReLU.
(a) Ablations (DQN)
(b) Ablations (DQN)
(c) Performance (PPO)
(d) Effective Rank (PPO)
@@
Figure 8: Baseline-normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), Tanh (HR) significantly outperforms Tanh. The Rational Activation (Delfosse et al., 2024) is comparable to ReLU within 40M frames. In (b), several ablations including layer normalization (Ba et al., 2016) and an increased dimension of $z_{t}\in\mathbb{R}^{1024}$ with a reduced learning rate $\alpha=5\mathrm{e}{-5}$ are shown. In b), Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product. Normalized performance (c) and effective rank (d) when training the PPO algorithm (Schulman et al., 2017). Note that no hyperparameter changes are done for the Hadamard representation.
@@
Examining further ablations in Fig. 8(a-b) shows that using another piece-wise linear function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product. We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$. Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
The Immunity of the ReLU
Interestingly, it seems that the ReLU activation’s performance is much less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations. As also indicated by Teney et al. (2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood. However, we have offered a partial explanation for this resilience of the ReLU in Section 4: ReLU saturation represents network pruning while hyperbolic tangent saturation leads to biases in subsequent layers.
@@
To evaluate results on a policy-based algorithm, additional experiments are run using the PPO algorithm (Schulman et al., 2017). For PPO, the internal architectural difference with DQN and PQN is that the final hidden layer $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients. After training for 40M frames, the performance and effective rank over time can be found in Fig. 8(c-d). Similarly to the results on DQN, the Tanh (HR) exhibits the highest effective rank and good performance. The Tanh (HR) and the ReLU seem to have consistent strong performance, whereas the pure hyperbolic tangent and sigmoid activations can be unreliable across algorithms. This is coherent with the fact that continuously differentiable activations are generally not favored over the ReLU (Teney et al., 2024).
@@
To test the performance of the Hadamard representation over a broader selection of environments, we evaluate the recent Parallelized Q-Network (PQN) algorithm in the nearly full Atari suite without hard-exploration environments, for a total of 51 games with 5 seeds per game.
@@
Figure 9: Median Human-Normalized scores on PQN in 51 Atari games for 5 seeds over 40M frames. Labels represent encoder activations. Employing a Hadamard representation provides a significant performance improvement over the recent PQN baseline (displayed as ReLU), the CReLU (Abbas et al., 2023) and ReDo ($\tau=0.025$) (Sokar et al., 2023). Notably, the improvement over a conventionally parameterized hyperbolic tangent activation is more than 100%.
@@
PQN is a vectorized-friendly version of DQN, reporting both a speed up from DQN as well as better convergence (Gallici et al., 2024). We run tests implementing the Hadamard representation in all hidden layers. The median human-normalized scores after training for 40M frames are shown in Fig. 9. We can clearly see a similar trend for the 51-game Atari suite. Using a Hadamard representation significantly increases the median performance as compared to the baseline PQN, the CReLU activation (Abbas et al., 2023), and the ReDo + ReLU algorithm (Sokar et al., 2023). Additionally, preliminary modified applications of ReDo to Tanh did not prove to be effective in PQN. However, we believe that future work in this direction would be worth pursuing.
6 Limitations
Using a Hadamard representation increases neural network parameters as it doubles the incoming weights connected to a hidden layer. However, in Fig. 9, a comparison against the CReLU is shown, which also nearly doubles the network’s parameters. Furthermore, recent work shows that simply scaling baselines in Atari often leads to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024). Another limitation is that, due to the vast amount of activation functions and the theoretical sound candidacy of the hyperbolic tangent, our research focused on the combination of hyperbolic tangents. Further research into Hadamard representations could find more novel activation combinations. A hyperparameter search will also give more insights into the strengths of the HR, as all baselines are specifically tuned for the ReLU. Finally, integration of a HR into more complex algorithms and architectures such as Rainbow (Hessel et al., 2018), Impala (Espeholt et al., 2018) or SPR Schwarzer et al. (2021) would be interesting.
Table 3: Parameter count across different activations
Activation Total Parameters
Tanh (Width2absent2$\cdot$ 2⋅ 2) 6,717,376
Tanh (HR) 3,373,382
CReLU 3,292,838
@@ 7 Conclusions and Discussion
This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem, but with more acute downstream effects. To alleviate this problem and make better use of these functions’ properties, a novel representation architecture called the Hadamard representation (HR) is proposed. The HR augments the hidden layers in a network by taking the Hadamard product with parallel, independently parameterized activation layers. We further analyzed and empirically showed that applying a Hadamard representation to hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness. In DQN, PQN and PPO, this approach significantly improved performance in Atari games compared to both standard representation parameterization and merely increasing the representation dimension. Future work could focus on further identifying the intricacies of the effects that different activation functions have on the resulting representation, in an attempt to push the potential of non piece-wise linear activations in reinforcement learning or even in a supervised learning setting (see Fig. 2). Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al. (2023) or in Delfosse et al. (2024) could be promising.